Two charges (βˆ’12.6 Γ— 10βˆ’12 𝐢𝐢 and βˆ’ 1.13 Γ— 10βˆ’13 𝐢𝐢) are separated by a distance of



3.75 πœ‡πœ‡π‘šπ‘š. Compute the force of attraction between them.

1
Expert's answer
2022-06-10T08:28:39-0400
"F=k\frac{q_1q_2}{r^2}"

"F=9*10^9\frac{\rm N\cdot m^2}{\rm C^2}*\frac{12.6*10^{-12}\: \rm C*1.13*10^{-13}\: C}{(3.75*10^{-6}\:\rm m)^2}\\
=9.11*10^{-4}\:\rm N"